OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x − y = 6 4x − 2y = 12 0 4 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system.

OpenStudy (anonymous):

Y don't u try it on ur own instead of me giving u all the answers!!!

OpenStudy (anonymous):

Once you know the metod it is very simple to solve

hero (hero):

y = 2x-6 y = (12-4x)/2 y = y 2(2x-6) = 12-4x 4x-12 = 12-4x 8x = 24 x = 3 y = 0